THE PURE ROTATIONAL SPECTRUM OF THE ZnSH RADICAL (X^{2}A′)

\underline{MATTHEW P. BUCCHINO} , GILLES R. ADANDE and LUCY M. ZIURYS, Department of Chemistry and Biochemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, Arizona 85721.

The pure rotational spectrum of the ZnSH radical (X^{2}A′) has been observed in the laboratory for the first time using millimeter/submillimeter direct-absorption methods and Fourier-transform microwave (FTMW) techniques in a frequency range of 4-400GHz. ZnSH was synthesized by reacting zinc vapor with H_{2}S under DC discharge in a Broida-type oven for the millimeter work; in the FTMW studies, the radical was created by discharge assisted laser ablation spectroscopy (DALAS) using 0.5% H_{2}S in Ar and a zinc rod. The K-ladder structure indicates C_{s} symmetry, and therefore a bent molecule. Spectra of multiple isotopologues have been recorded (^{64}ZnSH, ^{66}ZnSH, ^{68}ZnSH, and ^{64}ZnSD), from which an r_{0} structure has been determined. Each K-component consists of spin-rotation doublets with a splitting of 130-140MHz. Proton hyperfine structure was observed in the FTMW data. Rotational, spin-rotation, and hyperfine constants have been determined from a global fit to the data. Although ZnSH and ZnOH are isovalent, there appear to be subtle differences in bonding between the two species.